# Estimation of relative risk, odds ratio and their logarithms with guaranteed accuracy and controlled sample size ratio

### v3

@@ 1 Introduction
The rest of the paper is organized as follows. Section 2 describes the estimation procedure for RR, derives approximate expressions and bounds for the average sample sizes and estimation efficiency, and compares these with values obtained from Monte Carlo simulations. The estimation procedure and the results for LRR, OR and LOR are to some extent analogous, and are presented in Sects. 3, 4 and 5. Concluding remarks are given in Sect. 6.
@@ 2 Estimation of relative risk
The estimation procedure for RR with element sampling of the two populations is considered in Sect. 2.1. First, the general estimation approach is motivated and described (Sect. 2.1.1). The precise definition of the estimator is then completed (Sect. 2.1.2). Lastly, theoretical bounds are obtained for the average sample sizes and estimation efficiency, and these are compared with simulation results (Sects. 2.1.3 and 2.1.4).
Group sampling is addressed in Sect. 2.2. The estimation procedure used in this case is described (Sect. 2.2.1), and then approximations for the average number of groups and efficiency are obtained and compared with simulation results (Sects. 2.2.2 and 2.2.3).
@@ 2.1.1 Estimation procedure
$\frac{\operatorname{Var}\left[\hat{\theta}\right]}{\theta^{2}}=\frac{\operatorname{E}\left[\hat{\theta}^{2}\right]}{\theta^{2}}-\frac{\left(\operatorname{E}\left[\hat{\theta}\right]\right)^{2}}{\theta^{2}}=\frac{\operatorname{E}\left[\operatorname{E}\left[\hat{\theta}^{2}\mid s_{1},s_{2}\right]\right]}{\theta^{2}}-1<\operatorname{E}\left[e(s_{1},s_{2})\right]\leq A.$ (32)
@@ 2.1.2 Estimator parameters and approximate average sample sizes
There would be another pair of solutions where the square root in (34) has a negative sign, but that pair is not valid because it does not satisfy $s_{1}>\mu_{1}$, $s_{2}>\mu_{2}$ as required by (29). This is illustrated in Fig. 1, which makes it clear that there is only one solution pair in the valid range; and therefore it corresponds to the positive sign.
@@
As indicated in Sect. 2.1.1, the solutions (34) and (35) have to be rounded, because only integer numbers can be used as IBS parameters. Depending on their specific values it may be necessary to round both of them up, or it may be sufficient to round one up and the other down, if that satisfies (31). A simple criterion, which will be assumed in the rest of the paper, is as follows. First, randomly choose $(i,j)=(1,2)$ or $(2,1)$ with equal probability; then round $s_{i}$ up and $s_{j}$ down, and check if those values satisfy (31). If not, try rounding $s_{i}$ down and $s_{j}$ up. If not valid either, round both values up, which necessarily satisfies (31).
@@
Figure 2 represents these approximations using the example values (same as in Fig. 1) $A=0.05$, $\gamma=0.5$, $\delta_{1}=1$, $\delta_{2}=3$, and with $\mu_{1}$, $\mu_{2}$, $\mu_{12}$ given by (30) as specified for RR. The figure illustrates that the accuracy of the approximations depends on the curvature of $s_{1}$ and $s_{2}$ as functions of $X$ and $1/X$. In fact, for certain values of $A$, $\delta_{1}$ and $\delta_{2}$ the approximations (37) and (38) are exact. This will be analyzed in Sect. 2.1.3.
@@
The term $\xi$ models the average increase in $s_{1}$ and $s_{2}$ due to rounding. The impact of this on $\operatorname{E}[M_{i}+N_{1}]$, $i=1,2$ will usually be negligible, and thus $\xi$ could be taken as 0 with good approximation. However, choosing $\xi=1$ (together with appropriate values of $r_{1}$ and $r_{2}$) will be useful in Sect. 2.1.3 to obtain upper bounds on $\operatorname{E}[M_{i}+N_{i}]$.
@@ 2.1.3 Analysis of the approximation and bounds on average sample sizes
The error in approximating (34) and (35) by (37) and (38) vanishes when $s_{1}$ is an affine function of $X$ and $s_{2}$ is an affine function of $1/X$ (see Fig. 2). This happens when $A$, $r_{1}$ and $\xi$ satisfy a certain relationship, as discussed next. Expanding (66) and collecting terms, $D$ can be written as $d_{2}X^{2}+d_{1}X+d_{0}$ with
@@
Considering $r_{1}$ as given, this is a quadratic equation in $A$ with a single positive solution. Figure 3 shows the resulting curve for RR, i.e. for $\mu_{1}$, $\mu_{2}$, $\mu_{12}$ as in (30), with $\xi=1$. (The figure also contains curves for other estimators, to be presented in following sections.) For the pairs $(A,r_{1})$ in this curve the approximations (37) and (38) are exact. Note that $r_{1}$ is considered as a continuous variable for clarity of the representation, but only the points with integer $r_{1}$ (marked with dots in the graph) are feasible. Furthermore, it is seen from (76) that $c(A,r_{1},1)$ increases with $r_{1}$, and therefore the region above (below) the curve corresponds to $c(A,r_{1},1)$ positive (negative). Thus, for a given $A$, once the value of $r_{1}$ for which $c(A,r_{1},1)=0$ is known, rounding this up gives the minimum integer $r_{1}$ such that $c(A,r_{1},1)\geq 0$. That is, (81) is the integer on or immediately above the curve in Fig. 3, limited from below by $3$.
With the values selected for $\gamma$, $\delta_{1}$, $\delta_{2}$, $\epsilon_{1}$, $\epsilon_{2}$, $\xi$, $r_{1}$, $r_{2}$, the RR estimator is completely specified. The estimation procedure is summarized in Algorithm 1 (see Appendix A), together with a list of its properties, some of which will be derived in the remainder of this section. (The algorithm also includes the LRR case, to be presented in Sect. 3.)
@@
The following observations can be made from Fig. 4. For $r_{1}$ as in (81), simulated values are close to the theoretical bounds (and always below). When $A$ is not too large they are in fact very close, whereas for large $A$ the difference increases because $r_{1}$ is limited below by $3$. In addition, comparing with simulations for other values of $r_{1}$, (81) achieves average sample sizes equal or very close to their minima with respect to $r_{1}$. For $r_{1}$ given by (81), the ratio $\operatorname{E}[M_{1}+N_{1}]/\operatorname{E}[M_{2}+N_{2}]$ in the simulations is very close to $\lambda$ in all cases (for reference, note that the theoretical curves (79) and (80) are in the exact ratio $\lambda$); this will be analyzed with more detail later. The effect of $\phi$ on the normalized average sample sizes $\operatorname{E}[M_{i}+N_{i}]\phi$, $i=1,2$ is imperceptible (compare Figs. 5(b) and 5(c)). The parameter $\theta$ has an impact on the average sample sizes, but not on their ratio, which remains close to $\lambda$ (compare Figs. 4(a) and 4(b), or 4(c) and 5(a)). If both $\theta$ and $\lambda$ are replaced by their reciprocal values, $\operatorname{E}[M_{1}+N_{1}]$ and $\operatorname{E}[M_{2}+N_{2}]$ are simply swapped (see Figs. 4(c) and 5(b)). This is in agreement with (70) and (71). Lastly, for $A$ small the average sample sizes are approximately inversely proportional to this parameter. Again, this can be observed in (70) and (71), where for $A$ small the term $1/A$ dominates the other summands in the first factor.
@@
The theoretical curves in Fig. 4 have small jump discontinuities (see for example Fig. 4(c) near $A=0.07$), caused by the discrete character of $r_{1}$ and $r_{2}$. The jumps occur when the result of (81) changes by $1$. This effect is also present in the simulation results (bold dots), although less evident. Apart from this, for fixed $r_{1}$ the simulated results exhibit steep changes in certain locations, produced by the rounding applied to $s_{1}$ and $s_{2}$ (see for example the rightmost region for $\operatorname{E}[M_{2}+N_{2}]$ in Fig. 4(c) with $r_{1}=3$, which in that region corresponds to the bold dots). These are only observable in the simulation results, because the theoretical curves do not explicitly model the rounding of $s_{1}$ and $s_{2}$; and they are not discontinuities, but short sections where the slope is large in absolute value, as will be discussed later.
@@
Figure 6 compares the target $A$ with the relative MSE, $\operatorname{E}[(\hat{\theta}-\theta)^{2}]/\theta^{2}$, obtained from simulations in two specific cases, for $r_{1}$ given by (81). As seen, the relative MSE is always less than $A$, in accordance with (16). The difference between simulation and target is considerable when the latter is large. Again, this is a consequence of the limitation of $r_{1}$ in (81) to values not smaller than $3$. In addition, the difference increases slightly with $\phi$ (compare Figs. 6(a) and 6(b)). This is, at least in part, explained by the fact that the uniform bound (28) becomes less tight as $p_{1}$ or $p_{2}$ approach $1$, as can be seen by comparing it with (27).
@@
The relative MSE, like the average sample sizes, has steep changes near certain values of $A$, caused by rounding $s_{1}$ and $s_{2}$. The effect is noticeable in the rightmost region of Fig. 6, and is explained as follows. For large $A$, at least one of the two parameters $s_{1}$ and $s_{2}$ is small, according to (29) and (33), so as to make the relative MSE similar to (but smaller than) the target. Consequently that parameter, $s_{j}$, has a very narrow distribution before rounding, which implies that rounding it almost always produces the same integer value. This will typically be the next greater integer, because for small $s_{j}$ rounding causes a large variation in the error function, and thus rounding down is not likely to satisfy (31). In these conditions, if $A$ is increased by a small amount, the change in the distribution of $s_{j}$ before rounding, even if correspondingly small, can be sufficient to cause the rounded value to decrease by $1$, producing a significant change in $\operatorname{E}[(\hat{\theta}-\theta)^{2}]/\theta^{2}$, as well as in $\operatorname{E}[M_{j}+N_{j}]$.
@@
As a specific example, for $\lambda=1$, $\theta=16$ (Fig. 6(a)), consider $A=0.562$ (leftmost dot in the highest “plateau” of simulated values), for which (81) gives $r_{1}=3$. The simulation yields an average value of $1.856$ for $s_{2}$ before rounding, and a standard deviation of $0.082$. The value after rounding is $2$ with probability $0.962$ (and greater with probability $0.038$). The parameter $s_{1}$ before rounding takes larger values, with average $109.2$ and standard deviation $76.6$. Reducing $A$ to $0.501$ (next simulated value towards the left), which still corresponds to $r_{1}=3$, the average of $s_{2}$ before rounding becomes $2.087$, and the standard deviation is $0.094$. After rounding, $s_{2}$ now equals $3$ with probability $0.999$ (and is greater with probability $0.001$). Again, $s_{1}$ before rounding takes larger values, with mean and standard deviation $112.5$ and $77.9$ respectively. The change of the rounded value of $s_{2}$, from being $2$ to being $3$ with probabilities near $1$, causes the rightmost vertical gap in Fig. 6(a), from a relative MSE approximately equal to $0.50$ down to $0.34$. In contrast, for small $A$ both $s_{1}$ and $s_{2}$ are large and the effect of rounding is less marked, because, on one one hand, the distribution of each parameter is wider, with many possible rounded values; and, on the other hand, rounding a large value only causes a small variation in the relative MSE.
@@
The same behavior can be seen in the average sample sizes. Thus, in Fig. 4(b), which also corresponds to $\lambda=1$, $\theta=16$, the simulation results for $\operatorname{E}[M_{2}+N_{2}]$ with $r_{1}$ as in (81) (bold dots) show a vertical gap at the same horizontal position discussed in the preceding paragraph for the relative MSE. The effect is not discernible in the graph of $\operatorname{E}[M_{1}+N_{1}]$ because $s_{1}$ is large.
@@
It follows from the above that the effect of rounding $s_{1}$ and $s_{2}$ is not a discontinuity, but rather a short section with large slope in $\operatorname{E}[(\hat{\theta}-\theta)^{2}]/\theta^{2}$, or in $\operatorname{E}[M_{i}+N_{i}]$, as a function of $A$. Indeed, since (64) and (65) are continuous functions of $A$, the distributions of $s_{1}$ and $s_{2}$ also vary continuously with $A$, and so do the relative MSE and average sample sizes. This can be seen in Fig. 7(a), which is a detailed view of the rightmost part of Fig. 6(a) with smaller spacing along the horizontal axis, and in Fig. 7(b), which shows the corresponding values of $\operatorname{E}[M_{2}+N_{2}]$.
@@
The ratio $\operatorname{E}[M_{1}+N_{1}]/\operatorname{E}[M_{2}+N_{2}]$ obtained from simulation is represented in Fig. 8. As can be seen, it is in general close to $\lambda$, and very close for small or moderate values of $A$. The relatively larger deviations in the rightmost region of the graphs can again be attributed to the fact that for large $A$, since $r_{1}$ is limited from below by $3$ in (81), the value of $r_{1}$ does not result in $c(A,r_{1},\xi)$ close to 0. In any case, the deviation between the actual ratio and $\lambda$ is less than $11\%$ for all combinations of parameters shown in the figure, and much lower for $A$ small.

### v4

@@ 1 Introduction
The rest of the paper is organized as follows. Section 2 describes the estimation procedure for RR, derives approximate expressions and bounds for the average sample sizes and estimation efficiency, and compares these with values obtained from Monte Carlo simulations. The estimation procedure and the results for LRR, OR and LOR are to some extent analogous, and are presented in Sections 3, 4 and 5. Concluding remarks are given in Section 6.
@@ 2 Estimation of relative risk
The estimation procedure for RR with element sampling of the two populations is considered in Section 2.1. First, the general estimation approach is motivated and described (Section 2.1.1). The precise definition of the estimator is then completed (Section 2.1.2). Lastly, theoretical bounds are obtained for the average sample sizes and estimation efficiency, and these are compared with simulation results (Sections 2.1.3 and 2.1.4).
Group sampling is addressed in Section 2.2. The estimation procedure used in this case is described (Section 2.2.1), and then approximations for the average number of groups and efficiency are obtained and compared with simulation results (Sections 2.2.2 and 2.2.3).
@@ 2.1.1 Estimation procedure
$\frac{\operatorname{Var}\left[\hat{\theta}\right]}{\theta^{2}}=\frac{\operatorname{E}\left[\hat{\theta}^{2}\right]}{\theta^{2}}-\frac{\left(\operatorname{E}\left[\hat{\theta}\right]\right)^{2}}{\theta^{2}}=\frac{\operatorname{E}\left[\operatorname{E}\left[\hat{\theta}^{2}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right]}{\theta^{2}}-1<\operatorname{E}\left[e(s_{1},s_{2})\right]\leq A.$ (32)
@@ 2.1.2 Estimator parameters and approximate average sample sizes
There would be another pair of solutions where the square root in (34) has a negative sign, but that pair is not valid because it does not satisfy $s_{1}>\mu_{1}$, $s_{2}>\mu_{2}$ as required by (29). This is illustrated in Figure 1, which makes it clear that there is only one solution pair in the valid range; and therefore it corresponds to the positive sign.
@@
As indicated in Section 2.1.1, the solutions (34) and (35) have to be rounded, because only integer numbers can be used as IBS parameters. Depending on their specific values it may be necessary to round both of them up, or it may be sufficient to round one up and the other down, if that satisfies (31). A simple criterion, which will be assumed in the rest of the paper, is as follows. First, randomly choose $(i,j)=(1,2)$ or $(2,1)$ with equal probability; then round $s_{i}$ up and $s_{j}$ down, and check if those values satisfy (31). If not, try rounding $s_{i}$ down and $s_{j}$ up. If not valid either, round both values up, which necessarily satisfies (31).
@@
Figure 2 represents these approximations using the example values (same as in Figure 1) $A=0.05$, $\gamma=0.5$, $\delta_{1}=1$, $\delta_{2}=3$, and with $\mu_{1}$, $\mu_{2}$, $\mu_{12}$ given by (30) as specified for RR. The figure illustrates that the accuracy of the approximations depends on the curvature of $s_{1}$ and $s_{2}$ as functions of $X$ and $1/X$. In fact, for certain values of $A$, $\delta_{1}$ and $\delta_{2}$ the approximations (37) and (38) are exact. This will be analyzed in Section 2.1.3.
@@
The term $\xi$ models the average increase in $s_{1}$ and $s_{2}$ due to rounding. The impact of this on $\operatorname{E}[M_{i}+N_{1}]$, $i=1,2$ will usually be negligible, and thus $\xi$ could be taken as 0 with good approximation. However, choosing $\xi=1$ (together with appropriate values of $r_{1}$ and $r_{2}$) will be useful in Section 2.1.3 to obtain upper bounds on $\operatorname{E}[M_{i}+N_{i}]$.
@@ 2.1.3 Analysis of the approximation and bounds on average sample sizes
The error in approximating (34) and (35) by (37) and (38) vanishes when $s_{1}$ is an affine function of $X$ and $s_{2}$ is an affine function of $1/X$ (see Figure 2). This happens when $A$, $r_{1}$ and $\xi$ satisfy a certain relationship, as discussed next. Expanding (66) and collecting terms, $D$ can be written as $d_{2}X^{2}+d_{1}X+d_{0}$ with
@@
Considering $r_{1}$ as given, this is a quadratic equation in $A$ with a single positive solution. Figure 3 shows the resulting curve for RR, i.e. for $\mu_{1}$, $\mu_{2}$, $\mu_{12}$ as in (30), with $\xi=1$. (The figure also contains curves for other estimators, to be presented in following sections.) For the pairs $(A,r_{1})$ in this curve the approximations (37) and (38) are exact. Note that $r_{1}$ is considered as a continuous variable for clarity of the representation, but only the points with integer $r_{1}$ (marked with dots in the graph) are feasible. Furthermore, it is seen from (76) that $c(A,r_{1},1)$ increases with $r_{1}$, and therefore the region above (below) the curve corresponds to $c(A,r_{1},1)$ positive (negative). Thus, for a given $A$, once the value of $r_{1}$ for which $c(A,r_{1},1)=0$ is known, rounding this up gives the minimum integer $r_{1}$ such that $c(A,r_{1},1)\geq 0$. That is, (81) is the integer on or immediately above the curve in Figure 3, limited from below by $3$.
With the values selected for $\gamma$, $\delta_{1}$, $\delta_{2}$, $\epsilon_{1}$, $\epsilon_{2}$, $\xi$, $r_{1}$, $r_{2}$, the RR estimator is completely specified. The estimation procedure is summarized in Algorithm 1 (see Appendix A), together with a list of its properties, some of which will be derived in the remainder of this section. (The algorithm also includes the LRR case, to be presented in Section 3.)
@@
The following observations can be made from Figure 4. For $r_{1}$ as in (81), simulated values are close to the theoretical bounds (and always below). When $A$ is not too large they are in fact very close, whereas for large $A$ the difference increases because $r_{1}$ is limited below by $3$. In addition, comparing with simulations for other values of $r_{1}$, (81) achieves average sample sizes equal or very close to their minima with respect to $r_{1}$. For $r_{1}$ given by (81), the ratio $\operatorname{E}[M_{1}+N_{1}]/\operatorname{E}[M_{2}+N_{2}]$ in the simulations is very close to $\lambda$ in all cases (for reference, note that the theoretical curves (79) and (80) are in the exact ratio $\lambda$); this will be analyzed with more detail later. The effect of $\phi$ on the normalized average sample sizes $\operatorname{E}[M_{i}+N_{i}]\phi$, $i=1,2$ is imperceptible (compare Figures 5(b) and 5(c)). The parameter $\theta$ has an impact on the average sample sizes, but not on their ratio, which remains close to $\lambda$ (compare Figures 4(a) and 4(b), or 4(c) and 5(a)). If both $\theta$ and $\lambda$ are replaced by their reciprocal values, $\operatorname{E}[M_{1}+N_{1}]$ and $\operatorname{E}[M_{2}+N_{2}]$ are simply swapped (see Figures 4(c) and 5(b)). This is in agreement with (70) and (71). Lastly, for $A$ small the average sample sizes are approximately inversely proportional to this parameter. Again, this can be observed in (70) and (71), where for $A$ small the term $1/A$ dominates the other summands in the first factor.
@@
The theoretical curves in Figure 4 have small jump discontinuities (see for example Figure 4(c) near $A=0.07$), caused by the discrete character of $r_{1}$ and $r_{2}$. The jumps occur when the result of (81) changes by $1$. This effect is also present in the simulation results (bold dots), although less evident. Apart from this, for fixed $r_{1}$ the simulated results exhibit steep changes in certain locations, produced by the rounding applied to $s_{1}$ and $s_{2}$ (see for example the rightmost region for $\operatorname{E}[M_{2}+N_{2}]$ in Figure 4(c) with $r_{1}=3$, which in that region corresponds to the bold dots). These are only observable in the simulation results, because the theoretical curves do not explicitly model the rounding of $s_{1}$ and $s_{2}$; and they are not discontinuities, but short sections where the slope is large in absolute value, as will be discussed later.
@@
Figure 6 compares the target $A$ with the relative MSE, $\operatorname{E}[(\hat{\theta}-\theta)^{2}]/\theta^{2}$, obtained from simulations in two specific cases, for $r_{1}$ given by (81). As seen, the relative MSE is always less than $A$, in accordance with (16). The difference between simulation and target is considerable when the latter is large. Again, this is a consequence of the limitation of $r_{1}$ in (81) to values not smaller than $3$. In addition, the difference increases slightly with $\phi$ (compare Figures 6(a) and 6(b)). This is, at least in part, explained by the fact that the uniform bound (28) becomes less tight as $p_{1}$ or $p_{2}$ approach $1$, as can be seen by comparing it with (27).
@@
The relative MSE, like the average sample sizes, has steep changes near certain values of $A$, caused by rounding $s_{1}$ and $s_{2}$. The effect is noticeable in the rightmost region of Figure 6, and is explained as follows. For large $A$, at least one of the two parameters $s_{1}$ and $s_{2}$ is small, according to (29) and (33), so as to make the relative MSE similar to (but smaller than) the target. Consequently that parameter, $s_{j}$, has a very narrow distribution before rounding, which implies that rounding it almost always produces the same integer value. This will typically be the next greater integer, because for small $s_{j}$ rounding causes a large variation in the error function, and thus rounding down is not likely to satisfy (31). In these conditions, if $A$ is increased by a small amount, the change in the distribution of $s_{j}$ before rounding, even if correspondingly small, can be sufficient to cause the rounded value to decrease by $1$, producing a significant change in $\operatorname{E}[(\hat{\theta}-\theta)^{2}]/\theta^{2}$, as well as in $\operatorname{E}[M_{j}+N_{j}]$.
@@
As a specific example, for $\lambda=1$, $\theta=16$ (Figure 6(a)), consider $A=0.562$ (leftmost dot in the highest “plateau” of simulated values), for which (81) gives $r_{1}=3$. The simulation yields an average value of $1.856$ for $s_{2}$ before rounding, and a standard deviation of $0.082$. The value after rounding is $2$ with probability $0.962$ (and greater with probability $0.038$). The parameter $s_{1}$ before rounding takes larger values, with average $109.2$ and standard deviation $76.6$. Reducing $A$ to $0.501$ (next simulated value towards the left), which still corresponds to $r_{1}=3$, the average of $s_{2}$ before rounding becomes $2.087$, and the standard deviation is $0.094$. After rounding, $s_{2}$ now equals $3$ with probability $0.999$ (and is greater with probability $0.001$). Again, $s_{1}$ before rounding takes larger values, with mean and standard deviation $112.5$ and $77.9$ respectively. The change of the rounded value of $s_{2}$, from being $2$ to being $3$ with probabilities near $1$, causes the rightmost vertical gap in Figure 6(a), from a relative MSE approximately equal to $0.50$ down to $0.34$. In contrast, for small $A$ both $s_{1}$ and $s_{2}$ are large and the effect of rounding is less marked, because, on one hand, the distribution of each parameter is wider, with many possible rounded values; and, on the other hand, rounding a large value only causes a small variation in the relative MSE.
@@
The same behavior can be seen in the average sample sizes. Thus, in Figure 4(b), which also corresponds to $\lambda=1$, $\theta=16$, the simulation results for $\operatorname{E}[M_{2}+N_{2}]$ with $r_{1}$ as in (81) (bold dots) show a vertical gap at the same horizontal position discussed in the preceding paragraph for the relative MSE. The effect is not discernible in the graph of $\operatorname{E}[M_{1}+N_{1}]$ because $s_{1}$ is large.
@@
It follows from the above that the effect of rounding $s_{1}$ and $s_{2}$ is not a discontinuity, but rather a short section with large slope in $\operatorname{E}[(\hat{\theta}-\theta)^{2}]/\theta^{2}$, or in $\operatorname{E}[M_{i}+N_{i}]$, as a function of $A$. Indeed, since (64) and (65) are continuous functions of $A$, the distributions of $s_{1}$ and $s_{2}$ also vary continuously with $A$, and so do the relative MSE and average sample sizes. This can be seen in Figure 7(a), which is a detailed view of the rightmost part of Figure 6(a) with smaller spacing along the horizontal axis, and in Figure 7(b), which shows the corresponding values of $\operatorname{E}[M_{2}+N_{2}]$.
@@
The ratio $\operatorname{E}[M_{1}+N_{1}]/\operatorname{E}[M_{2}+N_{2}]$ obtained from simulation is represented in Figure 8. As can be seen, it is in general close to $\lambda$, and very close for small or moderate values of $A$. The relatively larger deviations in the rightmost region of the graphs can again be attributed to the fact that for large $A$, since $r_{1}$ is limited from below by $3$ in (81), the value of $r_{1}$ does not result in $c(A,r_{1},\xi)$ close to 0. In any case, the deviation between the actual ratio and $\lambda$ is less than $11\%$ for all combinations of parameters shown in the figure, and much lower for $A$ small.